OpenStudy (albert0898):

25) Three years ago Jane was one-half as old as she will be 2 years from now. What is Jane's present age?

OpenStudy (p0sitr0n):

janes age =x 2(x-3) = (x+2) 2x-6 = x+2 x = 8

OpenStudy (albert0898):

Positr0n is correct, I checked.

OpenStudy (p0sitr0n):

how can age be 2 if shes says 3 years before age =-1 right>

OpenStudy (albert0898):

Positr0n, can you just explain to me 'Three years ago Jane was one-half as old'? I under the three years ago part making it (x-3) but why'd you put times two right next to it. I don't get that.

OpenStudy (p0sitr0n):

beacause for the equality to be true, we need to take 2 the thing is it says in the past, as it is three years before so in the future, to make it true, we multiply think of x-3 as of a lower number to make the equality true, you need to increase it to be equal to a higher number thats it